# SuperDiff: Diffusion Models for Conditional Generation of Hypothetical New Families of Superconductors

### v2

@@ I Introduction
Currently, however, superconductivity in high $T_{c}$ superconductors is not very well understood. As a result, there exists no systematic method for searching for new high $T_{c}$ superconductors [1], and the most common method for searches for new high $T_{c}$ superconductors is essentially trial-and-error. For instance, the study in Ref. [2] surveyed approximately 1000 compounds over four years, of which they found only about $3\%$ to be superconducting. That study is a testament to the extreme inefficiency of finding new high $T_{c}$ superconductors through pure manual search.
@@
Understanding this, more recently, computational techniques have been applied to assist researchers in the search for new high $T_{c}$ superconductors. Specifically, a number of works have applied machine learning to this search for superconductors. Although serving as very valuable tools in many respects, most of these attempts [3, 4, 5], have been limited to classification and regression models, which only search through existing databases and are not able to generate any new compounds. Only recently, with deep generative models applied to superconductor discovery, have new hypothetical superconductors not found in most popular compound datasets been generated [6, 7, 8]. In Ref. [6], a Generative Adversarial Network (GAN) [9] was applied for unconditional high $T_{c}$ superconductor generation, and in Ref. [7], a Crystal Diffusion Variational Autoencoder (CDVAE) [10] was also applied for unconditional superconductor generation so that crystal structure could be accounted for. However, Ref. [7] was applied on a different dataset and focused on a different task of generating stoichiometric Bardeen–Cooper–Schrieffer (BCS) conventional superconductors [11] and so did not generate any superconductors with $T_{c}\gtrsim$ 20 K.
@@
New attempts at high $T_{c}$ superconductor discovery with generative models are not without limitations, however. Most notably, although past models have been able to successfully generate new superconductors within existing superconductor families, they have not been able to generate completely new families of superconductors, which would be particularly desirable. This is because they are only unconditional models, which learn only the training dataset distribution. As unconditional models, the generation process of these models cannot be controlled. In other words, past models lack conditioning functionality—a method for controlling the generation process, that, in this context, means giving an example superconductor, the reference compound, and having the model generate similar superconductors, ideally by interpolating between the example and what the model has learned from the training dataset. With conditioning, the possibility of generating new families of superconductors can be opened, and researchers can be given control over the generation process. This can be especially useful for researchers looking to find only specific types of superconductors or expand on their own new discoveries. Parallel to our work, Ref. [8] also applied a diffusion model for high $T_{c}$ superconductor discovery; however, their model was, like previous GANs, greatly limited by its lack of support for conditional generation with reference compounds—which is our main focus. Thus, their diffusion model shared with previous models the major limitation of being unable to generate any new families of superconductors—essentially, their work was only recreating the performance of the GAN in Ref. [6] but with a diffusion model instead and added $T_{c}$ label control only. Once again, we note that, in this work, we consider “conditioning” to mean conditioning the model on reference compounds only, as only this allows for the controlled generation of known and new families of superconductors. Moreover, Ref. [6] also struggled at generating unique (distinct from others in the given generated set) pnictides because of the small number of pnictides in SuperCon, the training dataset.
@@
To resolve these limitations, in this work, we implement a Denoising Diffusion Probabilistic Model (DDPM) [12, 13] for superconductor generation as our unconditional model and further implement conditioning with the Iterative Latent Variable Refinement (ILVR) [14] extension to DDPM, which allows for one-shot generation without additional training. With conditioning, we hope to be able to generate new families of superconductors for the first time, as identified by the clustering analysis proposed in Ref. [15], by experimenting with feeding the model different reference superconductors—this would mark a leap in the capabilities of computational searches for superconductors.
@@
We coin this first approach to conditionally generating new superconductors with reference compounds as “SuperDiff”. With SuperDiff, we aim to resolve the issues found in past works as a result of the small pnictide training dataset with the unconditional DDPM and, as our main focus, explore how the conditional DDPM can adapt to new information to generate completely new families of superconductors for the first time. This paper is structured as follows: in Section II, we give an overview of the underlying models behind SuperDiff and detail how they were applied; in Section III, we present the main results of the study, most notably a selection of new families of superconductors generated by SuperDiff—a first for the computational search for superconductors; and in Section IV, we summarize our contributions and most important findings.
@@ II Methodology
As stated in the introduction, we leverage the capabilities of Denoising Diffusion Probabilistic Models and Iterative Latent Variable Refinement to propose a method for conditionally generating new hypothetical superconductors. Here, we discuss the details of the creation of SuperDiff. In Section II.1, we discuss the sourcing and processing of superconductor data. We then provide a brief overview of the details of the underlying DDPM in Section II.2 and the Iterative Latent Variable Refinement method in Section II.3. Lastly, we discuss the techniques we use to evaluate the quality of SuperDiff outputs in Section II.4 and II.5.
@@ II.1 Data Processing
All data for the model was sourced from SuperCon[21], which is the largest database for superconducting materials. The dataset was processed by the steps in Ref. [6] and, like in previous studies [4, 15, 6], only the chemical composition data was used. Every compound from SuperCon was represented as a column vector for input into the model. As shown in Fig. 1, each compound was encoded as a $96\times 1$ column vector as $96$ is the maximum atomic number present in the dataset.
II.2 Denoising Diffusion Probabilistic Model
@@
where $\beta_{1},...,\beta_{T}$ is the variance schedule, $\mathbf{I}$ is the identity matrix, and $\mathbf{x}_{0}$ is dimensionally equivalent to latent variables $\mathbf{x}_{1},...,\mathbf{x}_{T}$ (all vectors in $\mathbb{R}^{96}$). In this work, we adopt the cosine variance schedule proposed in Ref. [22].
@@
In this work, we fix $\sigma_{t}^{2}=\beta_{t}$. Then, as shown in Ref. [12], by rewriting $\bm{\mu}_{\theta}$ as a linear combination of $\mathbf{x}_{t}$ and $\bm{\epsilon}_{\theta}$, a neural network that predicts $\bm{\epsilon}$ from $\mathbf{x}_{t}$ with input and output dimensions equal to that of the noise it predicts, the reverse process may be rewritten as:
@@
Four different versions of the DDPM were trained on SuperCon: one for cuprates, one for pnictides, one for others, and one for all classes (“everything”). The training datasets for each version of the DDPM were randomly split into training and validation sets in an approximately $95\%-5\%$ proportion. Training curves for all versions of the DDPM were able to converge and stabilize after around 50 epochs, and each version of the DDPM was trained for between 50 and 100 epochs, depending on the approximate lowest validation loss. For all versions of the DDPM, NAdam [25] was chosen as the optimizer, and provided satisfactory results. Moreover, like in Ref. [12], $T$ was set to 1000 and the U-Net [26] neural network architecture was used for $\bm{\epsilon}_{\theta}$ (for this work, a 1D U-Net was used as opposed to the 2D U-Net used for images).
@@ II.3 Conditioning
In Ref. [14], it was stated that the scale factor $N$ could be changed to control the amount of information brought from the reference to the generated output, where lower $N$ results in greater similarity between generated output and reference and higher $N$ results in only coarse information from the reference being brought by the model to the generated output. In our work, we found that $N>4$ resulted in large numbers of invalid compounds with negative amounts of elements. As a result, we used $N=2$ up to $N=4$, but we still found the conclusions made about changing $N$ in Ref. [14] applicable.
@@ II.4 Sampling
SuperDiff Version # Generated Inference Time (s)
Entirety of SuperCon 500,000 35,567
Cuprate 500,000 35,781
Pnictides 500,000 35,720
Others 500,000 35,679
Table 1: Inference metrics for the various versions of unconditional SuperDiff. These results were all achieved on a single consumer Nvidia RTX 3060 Ti GPU.
As mentioned previously, we trained four versions of the unconditional model, each of which was then copied and modified with ILVR conditioning to also create four versions of the conditional model. We thus have four versions of the unconditional DDPM (without ILVR), which we call “unconditional SuperDiff”, and four versions of the conditional DDPM (with ILVR), which we call “conditional SuperDiff”. On a single consumer Nvidia RTX 3060 Ti GPU, each version of SuperDiff was trained in under 2 hours, and we sampled 500,000 compounds from each of the four unconditional SuperDiff versions. Reasonably fast inference times were achieved, making SuperDiff a system that can be trained and used using resources at most universities and even consumers. Detailed inference metrics for unconditional SuperDiff are in Table 1. For conditional SuperDiff, we sampled varying amounts of compounds for different reference superconductors, and we discuss those results in Section III.5.
All sampled compounds were initially screened through various quality checks to ensure that all generated compounds were reasonably realistic. First, we obviously eliminated all generated compounds with negative amounts of elements. Note that we round all amounts of elements to two decimal places beforehand. Next, we eliminate compounds with either too few (only 1) or too many elements—for Cuprates, we limit outputs to compounds with a maximum of 7 elements, and for Pnictides and Others, we limit outputs to compounds with a maximum of 5 elements. After these basic checks, we removed duplicates and further evaluated compound validity with the charge neutrality and electronegativity checks from the SMACT package [27]. Finally, we ran formation energy prediction with ElemNet [28, 29]. We will discuss the performance of model generations against these checks later in Section III.
@@ II.5 Clustering
To identify if SuperDiff could generate new superconductor families, clustering analysis was performed. Clustering, which is an unsupervised machine-learning method used to find hidden patterns within data, was applied to the SuperCon database in Ref. [15], which established that these methods, when applied to superconductors, could exceed human performance in identifying different “families” of superconductors, which are represented as clusters. In this work, we use the clustering method for superconductors from Ref. [15] to evaluate generated outputs for new families. In Ref. [15], it was also found that, for superconductors, to visualize clustering results, the t-SNE method worked best. t-SNE is a non-linear dimensionality reduction technique that allows higher dimensional data ($96$-dimensional superconductor data points in this case) to be represented in 2D or 3D [30] (which do not have any physical meaning).
As discussed in the introduction, a major objective of this work was to generate new families of superconductors, as identified by the clustering model—that is, to generate new clusters of superconductors. This was something not accomplished by previous works, including the GAN in Ref. [6]. In order to achieve this goal, we experimented with the conditional model’s ability to interpolate between the reference compound and the training dataset. This idea of experimenting with a conditional DDPM’s ability to interpolate between the reference set and training set was proposed in Ref. [31] to attempt to achieve few-shot generation on image classes never seen during training. We attempt to do this with superconductors in this work. For instance, we experiment with conditioning the cuprate version of conditional SuperDiff on new, different reference cuprates not in the families of cuprate superconductors in the training dataset. We examine the model’s ability to generate new clusters or families of superconductors using information from the reference compound with this technique. We report our clustering results in Section III.4.
@@ III Results
In this section we report the performance of SuperDiff on various checks and discuss some interesting new findings. We first evaluate the performance of unconditional SuperDiff with the 500,000 compounds we generated for each of the four classes by performing various computational tests, which included some general compound checks in Sections III.1 and III.2 as well as checks for superconductivity in Section III.3. We use the same computational tests for unconditional SuperDiff as used for the GAN in Ref. [6] and are thus able to directly compare unconditional performance. Afterward, as our most notable results, we evaluate the performance of both the unconditional and conditional versions of SuperDiff on clustering in Section III.4 and manually identify and present some promising new families of superconductors generated by the conditional SuperDiff in Section III.5.
@@ III.1 Duplicates and Validity
SuperDiff Version Novel % Unique % # Valid
Entirety of SuperCon 100.00% 99.32% 79,828
Cuprate 100.00% 99.92% 10,971
Pnictides 100.00% 99.98% 2,184
Others 100.00% 99.39% 172,739
Table 2: The percentage of generated compounds that were novel (not in the training set) and unique (distinct from others in the given generated set) before SMACT [27] filters, and the number of generated compounds that were valid (passed SMACT filters) for the 500,000 compounds generated by each of the four versions of unconditional SuperDiff.
We notice that the novelty % and uniqueness % of generated results are all very high, which means that unconditional SuperDiff is able to successfully generate both diverse and novel compounds. Unconditional SuperDiff, here, outperforms the GAN in Ref. [6] in all metrics regarding generation novelty and uniqueness. Notably, unconditional SuperDiff maintains a very high uniqueness % for pnictides despite the small training set, something not accomplished by the Wasserstein GAN in Ref. [6]. This corroborates the observation of the superior ability of DDPMs to generate diverse results when compared to a GAN in other disciplines [16]. Lastly, although the SMACT check [27] results varied greatly between classes and the proportion of valid compounds for some classes was fairly low, the fast inference time justifies that SuperDiff is still able to generate valid compounds reasonably well for all classes.
Overall, these results indicate that all versions of unconditional SuperDiff are able to generate both novel and unique compounds—overcoming the past issues faced by Ref. [6]—as well as valid compounds. As conditional SuperDiff maintains much of the same components as the unconditional model, it was unsurprising that—in most cases—conditional SuperDiff was also able to generate novel, unique, and valid compounds; however, for conditional SuperDiff, these qualities were very much dependent on the reference compound—we still run these checks on all compounds generated by conditional SuperDiff and filter out invalid compounds.

### v3

# Diffusion Models for Conditional Generation of Hypothetical New Families of Superconductors

@@ I Introduction
Currently, however, superconductivity in high $T_{c}$ superconductors is not very well understood. As a result, there exists no systematic method for searching for new high $T_{c}$ superconductors [1], and the most common method for searches for new high $T_{c}$ superconductors is essentially trial-and-error. For instance, the study in Hosono et al. [2] surveyed approximately 1000 compounds over four years, of which they found only about $3\%$ to be superconducting. That study is a testament to the extreme inefficiency of finding new high $T_{c}$ superconductors through pure manual search.
@@
Understanding this, more recently, computational techniques have been applied to assist researchers in the search for new high $T_{c}$ superconductors. Specifically, a number of works have applied machine learning to this search for superconductors. Although serving as very valuable tools in many respects, most of these attempts [3, 4, 5], have been limited to classification and regression models, which only search through existing databases and are not able to generate any new compounds. Only recently, with deep generative models applied to superconductor discovery, have new hypothetical superconductors not found in most popular compound datasets been generated [6, 7, 8]. In Kim and Dordevic [6], a Generative Adversarial Network (GAN) [9] was applied for unconditional high $T_{c}$ superconductor generation, and in Wines et al. [7], a Crystal Diffusion Variational Autoencoder (CDVAE) [10] was also applied for unconditional superconductor generation so that crystal structure could be accounted for; however, that work used a different dataset and focused on the different task of generating stoichiometric Bardeen–Cooper–Schrieffer (BCS) conventional superconductors [11] and so did not generate any superconductors with $T_{c}\gtrsim$ 20 K.
@@
New attempts at high $T_{c}$ superconductor discovery with generative models are not without limitations, however. Most notably, although past models have been able to successfully generate new superconductors within existing superconductor families, they have not been able to generate completely new families of superconductors, which would be particularly desirable. This is because they are only unconditional models, which learn only the training dataset distribution. As unconditional models, the generation process of these models cannot be controlled. In other words, past models lack conditioning functionality—a method for controlling the generation process, that, in this context, means giving an example superconductor, the reference compound, and having the model generate similar superconductors, ideally by interpolating between the example and what the model has learned from the training dataset. With conditioning, the possibility of generating new families of superconductors can be opened, and researchers can be given control over the generation process. This can be especially useful for researchers looking to find only specific types of superconductors or expand on their own new discoveries. Parallel to our work, Zhong et al. [8] also applied a diffusion model for high $T_{c}$ superconductor discovery; however, their model was, like previous GANs, greatly limited by its lack of support for conditional generation with reference compounds—which is our main focus. Thus, their diffusion model shared with previous models the major limitation of being unable to generate any new families of superconductors—essentially, their work was only recreating the performance of the GAN in Kim and Dordevic [6] but with a diffusion model instead and added $T_{c}$ label control only. Once again, we note that, in this work, we consider “conditioning” to mean conditioning the model on reference compounds only, as only this allows for the controlled generation of known and new families of superconductors. Moreover, Kim and Dordevic [6] also struggled at generating unique (distinct from others in the given generated set) pnictides because of the small number of pnictides in SuperCon, the training dataset.
@@
To resolve these limitations, in this work, we implement a Denoising Diffusion Probabilistic Model (DDPM) [12, 13] for superconductor generation as our unconditional model and further implement conditioning with the Iterative Latent Variable Refinement (ILVR) [14] extension to DDPM, which allows for one-shot generation without additional training. With conditioning, we hope to be able to generate new families of superconductors for the first time, as identified by the clustering analysis proposed in Roter et al. [15], by experimenting with feeding the model different reference superconductors—this would mark a leap in the capabilities of computational searches for superconductors.
@@
We coin this first approach to conditionally generating new superconductors with reference compounds as “SuperDiff”. With SuperDiff, we aim to resolve the issues found in past works as a result of the small pnictide training dataset with the unconditional DDPM and, as our main focus, explore how the conditional DDPM can adapt to new information to generate completely new families of superconductors for the first time.
@@ II Methodology
As stated in the introduction, we leverage the capabilities of Denoising Diffusion Probabilistic Models and Iterative Latent Variable Refinement to propose a method for conditionally generating new hypothetical superconductors. Here, we discuss the details of the creation of SuperDiff by discussing the sourcing and processing of superconductor data, providing a brief overview of the details of the underlying DDPM and ILVR methods used, and discussing the techniques we use to evaluate the quality of SuperDiff outputs.
@@ II.1 Data Processing
All data for the model was sourced from SuperCon[21], which is the largest database for superconducting materials. The dataset was processed by the steps in Kim and Dordevic [6] and, like in previous studies [4, 15, 6], only the chemical composition data was used. Every compound from SuperCon was represented as a column vector for input into the model. As shown in Fig. 1, each compound was encoded as a $96\times 1$ column vector as $96$ is the maximum atomic number present in the dataset.
II.2 Denoising Diffusion Probabilistic Model
@@
where $\beta_{1},...,\beta_{T}$ is the variance schedule, $\mathbf{I}$ is the identity matrix, and $\mathbf{x}_{0}$ is dimensionally equivalent to latent variables $\mathbf{x}_{1},...,\mathbf{x}_{T}$ (all vectors in $\mathbb{R}^{96}$). In this work, we adopt the cosine variance schedule proposed in Nichol and Dhariwal [22].
@@
In this work, we fix $\sigma_{t}^{2}=\beta_{t}$. Then, as shown in Ho et al. [12], by rewriting $\bm{\mu}_{\theta}$ as a linear combination of $\mathbf{x}_{t}$ and $\bm{\epsilon}_{\theta}$, a neural network that predicts $\bm{\epsilon}$ from $\mathbf{x}_{t}$ with input and output dimensions equal to that of the noise it predicts, the reverse process may be rewritten as:
@@
Four different versions of the DDPM were trained on SuperCon: one for cuprates, one for pnictides, one for others, and one for all classes (“everything”). The training datasets for each version of the DDPM were randomly split into training and validation sets in an approximately $95\%-5\%$ proportion. Training curves for all versions of the DDPM were able to converge and stabilize after around 50 epochs, and each version of the DDPM was trained for between 50 and 100 epochs, depending on the approximate lowest validation loss. For all versions of the DDPM, NAdam [25] was chosen as the optimizer, and provided satisfactory results. Moreover, like in Ho et al. [12], $T$ was set to 1000 and the U-Net [26] neural network architecture was used for $\bm{\epsilon}_{\theta}$ (for this work, a 1D U-Net was used as opposed to the 2D U-Net used for images).
@@ II.3 Conditioning
In Choi et al. [14], it was stated that the scale factor $N$ could be changed to control the amount of information brought from the reference to the generated output, where lower $N$ results in greater similarity between generated output and reference and higher $N$ results in only coarse information from the reference being brought by the model to the generated output. In our work, we found that $N>4$ resulted in large numbers of invalid compounds with negative amounts of elements. As a result, we used $N=2$ up to $N=4$, but we still found the conclusions made about changing $N$ in Choi et al. [14] applicable.
@@ II.4 Sampling
SuperDiff Version Novel % Unique % # Valid Raw Output % True % Estimate Mean $T_{c}$ SD Max $T_{c}$
Everything 100.00% 99.32% 79,828 67.81% 62.22% $8.51\,\mathrm{K}$ $8.79\,\mathrm{K}$ $97.0\,\mathrm{K}$
Cuprates 100.00% 99.92% 10,971 64.53% 58.22% $64.68\,\mathrm{K}$ $14.61\,\mathrm{K}$ $110.5\,\mathrm{K}$
Pnictides 100.00% 99.98% 2,184 95.74% 96.39% $22.00\,\mathrm{K}$ $2.43\,\mathrm{K}$ $29.5\,\mathrm{K}$
Others 100.00% 99.39% 172,739 55.82% 47.56% $6.33\,\mathrm{K}$ $2.50\,\mathrm{K}$ $30.5\,\mathrm{K}$
Table 1: Summary of unconditional SuperDiff performance for the four versions we trained from the 500,000 compounds we sampled from each version. Shown are the percentage of generated compounds that were novel (not in the training set) and unique (distinct from others in the given generated set) before SMACT [27] filters, the number of generated compounds that were valid (passed SMACT filters), the percentages of generated compounds determined to be superconducting by the classification model along with the estimated true percentages according to Eq. 7, and summary $T_{c}$ statistics from the predictions by the regression model. We note that although the novelty percentage is 100.00%, this is due to rounding, and the model does, on extremely rare occasions, exactly reconstruct superconductors from the training dataset.
As mentioned previously, we trained four versions of the unconditional model, each of which was then copied and modified with ILVR conditioning to also create four versions of the conditional model. We thus have four versions of the unconditional DDPM (without ILVR), which we call “unconditional SuperDiff”, and four versions of the conditional DDPM (with ILVR), which we call “conditional SuperDiff”. On a single consumer Nvidia RTX 3060 Ti GPU, each version of SuperDiff was trained in under 2 hours, and we sampled 500,000 compounds from each of the four unconditional SuperDiff versions, which took less than 10 hours for each version. These relatively fast training and inference times make SuperDiff a system that can be trained and used using resources at most universities and even consumers. For conditional SuperDiff, we sampled varying amounts of compounds for different reference superconductors, and we discuss those results later.
All sampled compounds were initially screened through various quality checks to ensure that all generated compounds were reasonably realistic. First, we obviously eliminated all generated compounds with negative amounts of elements. Note that we round all amounts of elements to two decimal places beforehand. Next, we eliminate compounds with either too few (only 1) or too many elements—for Cuprates, we limit outputs to compounds with a maximum of 7 elements, and for Pnictides and Others, we limit outputs to compounds with a maximum of 5 elements. After these basic checks, we removed duplicates and further evaluated compound validity with the charge neutrality and electronegativity checks from the SMACT package [27]. Finally, we ran formation energy prediction with ElemNet [28, 29]. We will discuss the performance of model generations against these checks later.
@@ II.5 Clustering
To identify if SuperDiff could generate new superconductor families, clustering analysis was performed. Clustering, which is an unsupervised machine-learning method used to find hidden patterns within data, was applied to the SuperCon database in Roter et al. [15], which established that these methods, when applied to superconductors, could exceed human performance in identifying different “families” of superconductors, which are represented as clusters. In this work, we use the clustering method for superconductors from Roter et al. [15] to evaluate generated outputs for new families. In Roter et al. [15], it was also found that, for superconductors, to visualize clustering results, the t-SNE method worked best. t-SNE is a non-linear dimensionality reduction technique that allows higher dimensional data ($96$-dimensional superconductor data points in this case) to be represented in 2D or 3D [30] (which do not have any physical meaning).
As discussed in the introduction, a major objective of this work was to generate new families of superconductors, as identified by the clustering model—that is, to generate new clusters of superconductors. This was something not accomplished by previous works, including the GAN in Kim and Dordevic [6] and the diffusion model in Zhong et al. [8]. In order to achieve this goal, we experimented with the conditional model’s ability to interpolate between the reference compound and the training dataset. This idea of experimenting with a conditional DDPM’s ability to interpolate between the reference set and training set was proposed in Giannone et al. [31] to attempt to achieve few-shot generation on image classes never seen during training. We attempt to do this with superconductors in this work. For instance, we experiment with conditioning the cuprate version of conditional SuperDiff on new, different reference cuprates not in the families of cuprate superconductors in the training dataset. We examine the model’s ability to generate new clusters or families of superconductors using information from the reference compound with this technique, and we report our clustering results below.
@@ III Results
In this section we report the performance of SuperDiff on various checks and discuss some interesting new findings. We first evaluate the performance of unconditional SuperDiff with the 500,000 compounds we generated for each of the four classes by performing various computational tests, which included some general compound checks as well as checks for superconductivity. We use the same computational tests for unconditional SuperDiff as used for the GAN in Kim and Dordevic [6] and are thus able to directly compare unconditional performance. Afterward, as our most notable results, we evaluate the performance of both the unconditional and conditional versions of SuperDiff on clustering and manually identify and present some promising new families of superconductors generated by the conditional SuperDiff.
@@ III.1 Duplicates and Validity
We notice that the novelty % and uniqueness % of generated results are all very high, which means that unconditional SuperDiff is able to successfully generate both diverse and novel compounds. Unconditional SuperDiff, here, outperforms the GAN in Kim and Dordevic [6] in all metrics regarding generation novelty and uniqueness, and, similar to as proposed in their work, we also speculate that the high novelty percentage is due to the non-stoichiometric nature of the compounds we generate, which opens up a large composition space for the model. Notably, unconditional SuperDiff maintains a very high uniqueness % for pnictides despite the small training set, something not accomplished by the Wasserstein GAN in Kim and Dordevic [6]. This corroborates the observation of the superior ability of DDPMs to generate diverse results when compared to a GAN in other disciplines [16]. Lastly, although the SMACT check [27] results varied greatly between classes and the proportion of valid compounds for some classes was fairly low, the fast inference time justifies that SuperDiff is still able to generate valid compounds reasonably well for all classes.
Overall, these results indicate that all versions of unconditional SuperDiff are able to generate both novel and unique compounds—overcoming the past issues faced by Kim and Dordevic [6]—as well as valid compounds. As conditional SuperDiff maintains much of the same components as the unconditional model, it was unsurprising that—in most cases—conditional SuperDiff was also able to generate novel, unique, and valid compounds; however, for conditional SuperDiff, these qualities were very much dependent on the reference compound—we still run these checks on all compounds generated by conditional SuperDiff and filter out invalid compounds.